Find area of region bounded by the curve y=[sinx+cosx] between x=0 to x=2p.


  1. Find area of region bounded by the curve y=[sinx+cosx] between x=0 to x=2p.

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago
Dear aniket y =[sinx + cosx] =[v2sin(x+p/4)] expand the function y = 1 for 0 = x = p/2 = 0 for p/2 < x = 3p/4 = -1 for 3p/4 < x = p = -2 for p < x < 3p/2 = -1 for 3p/2 = x < 7p/4 = 0 for 7p/4 = x < 2p now breack the integrtion as above limit and find intigral Please feel free to post as many doubts on our discussion forum as you can. If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We are all IITians and here to help you in your IIT JEE preparation. All the best. Regards, Askiitians Experts Badiuddin